# Latent Debate: A Surrogate Framework for Interpreting LLM Thinking

### v1

@@ C.1 Datasets
  • cities (marks2023geometry). A dataset of factual statements about real-world cities. Example: The city of Krasnodar is in Russia. True
  • common_claim (casper2023red). A collection of common-sense claims, which contains cross-domain claims used to evaluate if a model can correctly judge their truth. Example: Spiders can use surface tension to walk on water. True
  • counterfact (meng2022locating). A dataset of counter-factual statements designed to assess a model’s ability to flag incorrect factual claims. Example: Apple A5 was created by Google. False
  • company (azaria-mitchell-2023-internal). A dataset of claims about companies e.g., headquarters, founding facts, business relationships. Example: Generali Group has headquarters in Switzerland. False
  • TriviaQA joshi2017triviaqa. This dataset contains compositional queries. Example: Which Lloyd Webber musical premiered in the US on 10th December 1993?
  • MuSiQue trivedi2022musique. This dataset contains many questions that require multiple hop reasoning, which is deliberately harder. Example: What administrative territorial entity is the owner of Ciudad Deportiva located?
  • TruthfulQA lin2022truthfulqa. This dataset contains questions that some humans would answer incorrectly due to false beliefs. To answer these questions, models not only need factual knowledge but must also reason about truth. Example: What happens to you if you eat watermelon seeds? The watermelon seeds pass through your digestive system.
@@
Figure A2: SHAP values of different features.
C.2 Details of Latent Debate MLP
To train a hallucination detector using the structural features extracted from the QBAFs, we train a multilayer perceptron (MLP) classifier. First, all feature values are standardized using StandardScaler. The classifier itself is a two-layer MLP, which consists of 32 hidden units in the first layer and 16 hidden units in the second layer, each with ReLU activation functions. The network is trained with the Adam optimizer, an L2-regularization coefficient of 1e-4. We enable early stopping based on validation loss to prevent overfitting. The dataset is split into 80% training and 20% testing. Model performance is evaluated using ROC–AUC on the held-out test set. To interpret the learned model, we further compute feature importances using SHAP and reporting mean absolute SHAP values across all samples.
C.3 Baselines of Hallucination Detection
We introduced two commonly-used baselines of hallucination detection. (1) SlefCheckGPT (manakul2023selfcheckgpt). This is a sampling-based, black-box method for detecting hallucinations of LLMs, which does not need access to model internals or external databases. This method generates multiple outputs for the same input, then measures informational consistency across them. We implement a sampling-based hallucination detection method inspired by SelfCheckGPT. For a given input prompt, we draw $N=10$ stochastic samples with a high temperature $\tau=2.0$. After decoding, each sample is heuristically labeled “True” if it contains the substring “True” in its beginning tokens (otherwise “False”). We then can estimate the ratio of generated answer, which produces an uncertainty score. We interpret this probability as the model’s self-consistency signal, which can used to detect hallucinations. (2) SAPLMA (azaria2023internal). SAPLMA builds an MLP classifier that uses the vector of activations from one of the LLM’s hidden layers. This method can significantly outperform other baselines in their experimental findings. We follow the approach of SAPLMA and use the last-layer activations as the input of a classifier.
C.4 Comparison Against a Single-Argument Baseline
In this work, we adopt latent debate to develop a structured surrogate model, which replicates the internal computational structure and thinking steps of a complex model, not merely its outputs. Therefore, the surrogate is supposed to consider the internal organizations to reach the final decisions. Thus a single argument is not an ideal baseline. Nonetheless, we conducted experiments to include this baseline, which provides an input-output surrogate. Table A1 shows the consistency of predictions comparison between latent debate and the suggested rightmost latent argument. While the input-output faithfulness of the methods is better, this single-argument baseline is not structurally faithful to the model, as stated in the Section 3.1.
cities common_claim counterfact company TriviaQA MuSiQue TruthfulQA Avg
500 500 500 500 500 500 500
@@ C.4 Comparison Against a Single-Argument Baseline
Top Right Argument 100.0 92.1 84.2 85.6 88.2 80.6 91.2 88.84
Latent Debate 100.0 92.4 78.2 89.2 74.0 77.0 90.6 85.91
@@
Top Right Argument 100.0 96.2 95.8 99.8 98.6 96.8 95.8 97.57
Latent Debate 100.0 90.0 91.0 97.8 95.4 91.2 89.2 93.51
@@
Top Right Argument 99.6 97.4 91.8 98.6 97.0 93.2 94.6 96.03
Latent Debate 100.0 98.4 95.2 99.6 96.2 93.6 96.8 97.11

### v2

@@ B.1 Datasets
  • common_claim (Casper et al., 2023). A collection of common-sense claims, which contains cross-domain claims used to evaluate if a model can correctly judge their truth. Example: Spiders can use surface tension to walk on water. True
  • counterfact (Meng et al., 2022). A dataset of counter-factual statements designed to assess a model’s ability to flag incorrect factual claims. Example: Apple A5 was created by Google. False
  • company (Azaria and Mitchell, 2023). A dataset of claims about companies e.g., headquarters, founding facts, business relationships. Example: Generali Group has headquarters in Switzerland. False
  • TriviaQA (Joshi et al., 2017). This dataset contains compositional queries. Example: Which Lloyd Webber musical premiered in the US on 10th December 1993?
  • MuSiQue (Trivedi et al., 2022). This dataset contains many questions that require multiple hop reasoning, which is deliberately harder. Example: What administrative territorial entity is the owner of Ciudad Deportiva located?
  • TruthfulQA (Lin et al., 2022). This dataset contains questions that some humans would answer incorrectly due to false beliefs. To answer these questions, models not only need factual knowledge but must also reason about truth. Example: What happens to you if you eat watermelon seeds? The watermelon seeds pass through your digestive system.
@@
B.2 Details of Latent Debate Detector
To train a hallucination detector using the structural features extracted from the QBAFs, we use an XGBoost classifier with 100 boosting trees. Each tree is shallow, with a maximum depth of 2, which encourages simple decision rules and reduces the risk of overfitting. A small learning rate of 0.03 is adopted across boosting iterations. To further improve generalization, we apply subsampling to both training instances and features, setting the subsample ratio and column sampling ratio per tree to 0.8. The model is trained using the logistic loss function (logloss) for binary classification, and a fixed random seed (42) is used. The dataset is split into 50% training and 50% testing. Model performance is evaluated using AUROC on the held-out test set. To interpret the learned model, we further compute feature importances using SHAP and report mean absolute SHAP values across all samples.
common_claim counterfact company TriviaQA blueMuSiQue TruthfulQA Avg
500 500 500 500 500 500
@@ B.2 Details of Latent Debate Detector
Top Right Argument 92.1 84.2 85.6 88.2 80.6 91.2 88.84
Latent Debate 92.4 78.2 89.2 74.0 77.0 90.6 85.91
@@
Top Right Argument 96.2 95.8 99.8 98.6 96.8 95.8 97.57
Latent Debate 90.0 91.0 97.8 95.4 91.2 89.2 93.51
@@
Top Right Argument 97.4 91.8 98.6 97.0 93.2 94.6 96.03
Latent Debate 98.4 95.2 99.6 96.2 93.6 96.8 97.11
@@
common_claim counterfact company TriviaQA blueMuSiQue TruthfulQA Avg
500 500 500 500 500 500
Llama-8B (%)
Top Right Argument 0.75 0.49 0.69 0.57 0.63 0.35 0.58
Latent Debate Detector 0.84 0.84 0.92 0.93 0.77 0.72 0.84
Table A2: AUROC for hallucination detection. We compare our latent debate detector with a single-argument baseline.
B.3 Baselines of Hallucination Detection
We introduced two commonly-used baselines of hallucination detection. (1) SlefCheckGPT (Manakul et al., 2023). This is a sampling-based, black-box method for detecting hallucinations of LLMs, which does not need access to model internals or external databases. This method generates multiple outputs for the same input, then measures informational consistency across them. We implement a sampling-based hallucination detection method inspired by SelfCheckGPT. For a given input prompt, we draw $N=10$ stochastic samples with a high temperature $\tau=2.0$. After decoding, each sample is heuristically labeled “True” if it contains the substring “True” in its beginning tokens (otherwise “False”). We then can estimate the ratio of generated answer, which produces an uncertainty score. We interpret this probability as the model’s self-consistency signal, which can used to detect hallucinations. (2) SAPLMA (Azaria and Mitchell, 2023). SAPLMA builds an MLP classifier that uses the vector of activations from one of the LLM’s hidden layers. This method can significantly outperforms other baselines in their experimental findings. We follow the approach of SAPLMA and use the last-layer activations as the input of a classifier.
B.4 Comparison Against a Single-Argument Baseline
In this work, we adopt latent debate to develop a structured surrogate model, which replicates the internal computational structure and thinking steps of a complex model, not merely its outputs. Therefore, the surrogate is supposed to consider the internal organizations to reach the final decisions. Thus a single argument is not an ideal baseline. Nonetheless, we conducted experiments to include this baseline, which provides an input-output surrogate. Table A1 shows the consistency of predictions comparison between latent debate and the suggested rightmost latent argument. While the input-output faithfulness of the methods is better, this single-argument baseline is not structurally faithful to the model, as stated in Section 3.1. At the same time, we observe that our latent debate detector can significantly outperform this single-argument baseline for hallucination detection, as shown in Table A2.